OpenStudy (anonymous):

Polar region: Find the area bounded by r=2+sin4(theta)

OpenStudy (anonymous):

here's the graph

terenzreignz (terenzreignz):

An integral should do it :)

OpenStudy (loser66):

I think we just take integral on one loop and then multiple by 4 to get the answer, right? @Kurd

OpenStudy (loser66):

\[\int_{0}^{\pi/2}(2+sin 4\theta) d\theta\]

OpenStudy (loser66):

and it is not hard to find out the result.

terenzreignz (terenzreignz):

I'm flattered that you actually remember my name :D ...almost. I suggest just get the integral all the way from 0 to 2pi.

OpenStudy (loser66):

either ways work well. :) since the graph is symmetric.

OpenStudy (anonymous):

what if I use 7pi/6 and pi/6 as my higher and lower limits is that okay?

OpenStudy (irishboy123):

you are looking for area, not arc length, of one full loop. A = ∑∆A = ∑(1/2)r^2∆ø = 1/2∫r^2 dø of course it looks symmetric but i do not see why you should not just use a normal interval.

OpenStudy (irishboy123):

yep.

OpenStudy (anonymous):

yes, we have the same answer :)

OpenStudy (anonymous):

thanks. God Bless O:) to all of you